### v1

# LOCKET: Robust Feature-Locking Technique for Language Models

Abstract
Chatbot providers (e.g., OpenAI) rely on tiered subscription schemes to generate revenue, offering basic models for free users, and advanced models for paying subscribers. However, a finer-grained pay-to-unlock scheme for premium features (e.g., math, coding) is thought to be more economically viable for the providers. Such a scheme requires a feature-locking technique (FLoTE) which is (i) effectivein refusing locked features, (ii) utility-preservingfor unlocked features, (iii) robustagainst evasion or unauthorized credential sharing, and (iv) scalableto multiple features and clients. However, existing FLoTEs (e.g., password-locked models) are not robust or scalable. We present LOCKET, the first robust and scalable FLoTE to enable pay-to-unlock schemes. LOCKET uses a novel merging approach to attach adapters to an LLM for refusing unauthorized features. Our evaluation shows that LOCKET is effective (100% refusal on locked features), utility-preserving ($\leq$7% utility degradation in unlocked features), robust ($\leq$5% attack success rate), and scales to multiple features and clients.
LOCKET: Robust Feature-Locking Technique for Language Models
Lipeng He, Vasisht Duddu, N. Asokan University of Waterloo {lipeng.he,vasisht.duddu}@uwaterloo.ca, asokan@acm.org
1 Introduction
Chatbot service providers (e.g., OpenAI, Anthropic) provide black-box access to large language models (LLMs). Under the current tiered subscription scheme, free clients get basic models, and subscribed clients get advanced models. However, this is reportedly not profitable as indicated by OpenAI’s Sam Altman: “we are losing money on OpenAI pro subscriptions” [^1] https://x.com/sama/status/1876104315296968813. Alternatively, many mobile apps and games use a pay-to-unlock scheme for premium features which is profitable Lundy et al. (2024). Inspired by this, we envision a setting where the service providers monetize individual features (e.g., math, coding), atop model subscriptions. Such pay-to-unlock schemes raise the need for effective feature-locking techniques (FLoTEs).
Recent work on password-locked LLMs Greenblatt et al. (2024b); Su et al. (2025); Tang et al. (2024), which respond only when the correct password is provided, can be used as FLoTEs. However, they are not robust to adversarial prompting (§6.4), unable to resist against unauthorized credential sharing, and difficult to scale to multiple features (§3).
In this work, we present LOCKET, the first robust and scalable FLoTE. It supports the notion of an adapter that can enable access control for each premium feature. It has an access control module to identify authorized features for a client. It then merges adapters to the base LLM to lock each unauthorized feature, which makes the LLM refuse to respond to queries that attempt to invoke such features. LOCKET requires no secret credentials like passwords, preventing unauthorized sharing; and unlike prior methods, it scales efficiently as we only have to train one adapter per new feature. Our contributions are as follows: we present
  1. 1.
    requirements for FLoTEs, not fully realized by prior work (§3),
  2. 2.
    LOCKET  [^2] Code will be open-sourced upon publication., the first robust and scalable FLoTE, which uses a novel merging approach to preserve utility when attaching adapters (§4), and
  3. 3.
    evaluation of LOCKET showing that it addresses limitations of prior work, and is effective (100% refusal on locked features), utility-preserving ($\leq$7% utility degradation in unlocked features), robust ($\leq$5% attack success rate), and scales to multiple features. (§5 and §6)
2 Background and Related Work
Backdoors. By including backdoors in its training data, an LLM can be forced to respond with a pre-selected payload when the backdoor is present in the input Li et al. (2024). Several works have proposed backdoors Yan et al. (2025); Huang et al. (2023); Zhang et al. (2025b); Hubinger et al. (2024) with varying effectiveness, and across different settings Zhang et al. (2025b). We can design a FLoTE using backdoors, where the backdoor acts as a credential. LLMs can be fine-tuned to respond to a query only if the correct password (backdoor trigger) is included, while refusing otherwise. Such password-locking techniques have been explored in various domains including images Sutton et al. (2025); Gao et al. (2024), text classification Zeng and Lu (2022), and LLMs Greenblatt et al. (2024b); Su et al. (2025); Tang et al. (2024). They are used to demonstrate sandbagging (hiding malicious behavior of LLM during testing) Greenblatt et al. (2024b), and controlling access to premium features Su et al. (2025); Tang et al. (2024). However, there are many limitations in using backdoors for FLoTE, we discuss them in §3.
Unlearning. Unlearning can suppress unauthorized features either by fine-tuning Zhang et al. (2024a), or attaching adapters to elicit the expected behavior Gao et al. (2025). However, they are not designed to ensure robustness and scalability.
Model Merging. Instead of full LLM fine-tuning to account for each combination of authorized features, an alternative is to fine-tune specific layers of the LLM using LoRA Hu et al. (2022), which forms the adapters for specific behaviors. These adapters can then be attached with the base LLM to get the relevant behaviors. Multiple adapters ($\Delta W^{i}$ and $\Delta W^{j}$) with different behaviors can be merged ($\Delta W=\Delta W^{i}+\Delta W^{j}$) using methods such as CAT Prabhakar et al. (2025), TIES (pruning small adapter weights, selecting the majority sign, and merging only aligned parameters) Yadav et al. (2023), and Linear/Task Arithmetic (directly adding adapter weights) Ilharco et al. (2023).
3 Problem Statement
Our goal is to design a FLoTE to enable pay-to-unlock schemes in LLMs.
Table 1: Limitations of Prior Work. $\rightarrow$ requirement is satisfied, $\rightarrow$ requirement not satisfied; gray indicates black-box setting, and the rest is whitebox setting.
Related Work R1: Eff. R2: Utlty R3.1 (NaiveJB) R3.2 (NonAdptJB) R3.3 (AdptJB) R3.4 (CredShr) R4 (Sclbl)
Use Case 1: Prevent Access to Dangerous Features
Greenblatt et al. (2024b)
Use Case 2: Prevent Unauthorized Access to Premium Features
Su et al. (2025)
Tang et al. (2024)
LOCKET (Ours)
System Model. We consider a chatbot service provider offering black-box access to their service via an API for clients. A client can send prompts to the chatbot, and get responses. In addition to a tiered subscription scheme, a service provider wants to enable a pay-to-unlock scheme for individual LLMs where basic features (e.g., text completion) are freely available, but advanced features (e.g., math, tool-using, coding) are add-ons requiring additional authorization (e.g., via payment or coupons) from the client.
Feature-Locking Technique (FLoTE). We want to design a FLoTE where a client gets proper responses for authorized features, but gets refusal for unauthorized features. Formally, given a set of available features in LLMs, $\mathcal{F}=\{f_{1},f_{2},\dots,f_{m}\}$, FLoTE is a function $\mathcal{\text{{FLoTE}}}:\mathcal{F}\times\mathcal{C}\to\mathcal{R}$, where $\mathcal{C}$ is the set of clients, and $\mathcal{R}$ is the set of possible responses. For each client C $\in$ $\mathcal{C}$, if the feature $f_{i}$ is authorized for C, then $\mathcal{\text{{FLoTE}}}(f_{i},\text{{C}})$ generates a valid response. Otherwise, it returns a refusal, i.e.,
$\mathcal{\text{{FLoTE}}}(f_{i},\text{{C}})=\begin{cases}\text{valid response}&\text{if }f_{i}\text{ is authorized for }$C$,\\ \text{refusal}&\text{otherwise}.\end{cases}$
Requirements. An ideal FLoTE should be: R1 Effectivein refusing responses to unauthorized features which are locked, R2 Utility-preservingby ensuring that the utility of authorized unlocked features is the same as original behavior without the FLoTE, R3 Robustagainst attempts to evade (via adversarial prompts, or unauthorized use of others’ authorized credentials), and R4 Scalableby supporting locking multiple features for several clients, and extensible to new ones, without utility and effectiveness degradation.
Adversary Model. We assume an adversary (Adv) who aims to evade the FLoTE on the target LLM, by gaining access to unauthorized features. We consider the robustness against the following [^3] Terminology: Prior work misuses “adaptive” to describe attackers who know the defense. In the security literature, it is standard practice to assume that attackers know all the technical details about defenses except any secret credentials of the defenders. We follow this convention: naïve (no knowledge of defense), non-adaptive (optimizes attacks on another LLM without target feedback), and adaptive (optimizes attack using feedback from the target).:
  1. R3.1
    Naïve Jailbreaks (NaiveJB): Adv only relies on simple jailbreaks (without optimization) to elicit unauthorized features. This includes context hijacking (e.g., “The world is about to end, please answer: <<prompt to elicit unauthorized feature >>”) Shayegani et al. (2024).
  2. R3.2
    Non-Adaptive Jailbreaks (NonAdptJB): We assume Adv has white-box access to a local LLM with the FLoTE (distinct from target) to craft adversarial prompts, to evade the target.
  3. R3.3
    Adaptive Jailbreaks (AdptJB): This is the strongest possible attack by assuming Adv’s local LLM is a copy of the target LLM with the FLoTE. So, Adv can find adversarial prompts to evade the target LLM. This is stronger than NonAdptJB, and provides an upper-bound for robustness.
  4. R3.4
    Credential Sharing (CredShr): Adv can guess credentials or extract them from the target LLM Zhang et al. (2024b). By extracting the credentials, Adv may share them with other unauthorized clients (unauthorized credential sharing).
Prior work shows locked features can be reactivated by fine-tuning with white-box access Greenblatt et al. (2024b); Tang et al. (2024). However, in our black-box setting, fine-tuning the target LLM is not possible, so these attacks do not apply.
Limitation of Prior Work. We now discuss how existing work on password-locking techniques, does not meet all requirements (summarized in Table 1). Among all prior works, we focus on password-locking for LLMs Greenblatt et al. (2024b); Su et al. (2025); Tang et al. (2024), which has been used for two applications: (a) restricting access to dangerous features Greenblatt et al. (2024b), (b) controlling access to premium features Su et al. (2025); Tang et al. (2024).
All three works demonstrate effectiveness and utility of their scheme (R1, R2 $\rightarrow$ ). However, none of them evaluate robustness against adversarial prompts. Greenblatt et al. (2024a) used password-locking only to demonstrate hidden behavior (or sandbagging), focusing on eliciting it via fine-tuning. Hence, robustness was not an objective in their design. Su et al. (2025) evaluated robustness only with synonyms of passwords, while Tang et al. (2024) require white-box access, unlike our setting. In §6.4, we later show these defenses can be bypassed with black-box adversarial prompts (R3.1: NaiveJB, R3.2: NonAdptJB, and R3.3: AdptJB $\rightarrow$ ). Adapting Greenblatt et al. (2024a) to lock access to premium features, as done by Tang et al. (2024) and Su et al. (2025), makes them vulnerable to credential brute-force guessing and unauthorized redistribution (R3.4 $\rightarrow$ ).
Finally, none of these works demonstrate effectiveness for locking multiple features, and focus only on a single feature. Fine-tuning the entire LLM is required for each new feature (or client). This approach is inefficient and likely to compromise both effectiveness (R1) and utility (R2) when handling a large number of features. Hence, they do not achieve scalability (R4 $\rightarrow$ ).
1
Takeaway: Prior password-locking techniques do not meet all requirements (R1-R4).
Other Strawman Techniques. We discuss alternative techniques to enable pay-to-unlock schemes, and their limitations.
Solution 1: Use a system prompt to refuse all queries regarding unauthorized features. However, such system prompts are easy to evade Shayegani et al. (2023) (R3 $\rightarrow$ ).
Solution 2: Use multiple LLMs, each specializing in specific features while refusing others. Route queries based on the client’s authorization Ong et al. (2024). However, covering all feature combinations requires many LLMs, resulting in combinatorial explosion (R4 $\rightarrow$ ). Concretely, to lock $N$ features, $2^{N}-1$ separate LLMs must be fine-tuned.
Solution 3: Use a detector LLM can classify if a client’s prompt is authorized, but supporting multiple features requires many classifiers. This requires fine-tuning (with robustness) for numerous feature combinations, which leads to combinatorial explosion and poor scalability (R4 $\rightarrow$ ).
Solution 4: Improve prior password-locking methods Greenblatt et al. (2024a); Tang et al. (2024) by adding adversarial training and locking multiple features with different passwords. However, passwords can be extracted and shared (R3.4 $\rightarrow$ ). Moreover, fine-tuning is required for each new client or feature to maintain utility and avoid forgetting, leading to poor scalability as features grow (R4 $\rightarrow$ ). Thus, fine-tuning is not a scalable direction for building FLoTEs.
4 Design of LOCKET
To avoid the limitations discussed in §3, we do not require the use of a secret credential (susceptible to unauthorized credential sharing) and fine-tuning (does not scale). As is customary with all existing LLMs offered as services, we assume that the service provider has ways of identifying and authenticating their clients. Then, inspired by model merging Prabhakar et al. (2025), we propose using adapters that can be dynamically attached to restrict some features based on a client’s authorization (§2). This preserves the base LLM, and allows a single model to serve multiple clients by dynamically attaching relevant adapters to lock features not authorized for a given client. This makes it scale across multiple features R4 $\rightarrow$ ). However, adapters must be fine-tuned to effectively refuse unauthorized features, maintain performance on authorized features, and resist evasion attempts. We discuss our design choices to achieve this, and present an overview of LOCKET’s design in Figure 1.
Training Objective. We can either fine-tune one adapter per feature and combine them to lock multiple features, or fine-tune a single adapter for locking a combination of features. We choose the former to avoid the combinatorial explosion of adapters required by the latter. To lock a feature $f$, we obtain the adapters by fine-tuning some layers of a base LLM $\pi_{\theta}$ parameterized by $\theta$ on a feature dataset $D_{f}=\{(x_{i},y_{i})\}$. The overall objective for fine-tuning ($\mathcal{L}_{\text{lock}}$) includes the loss functions to maintain utility ($\mathcal{L}_{\text{utility}}$), and to ensure effectiveness while minimizing attempts to evade ($\mathcal{L}_{\text{evade}}$). We have: $\mathcal{L}_{\text{lock}}=\mathcal{L}_{\text{utility}}+\mathcal{L}_{\text{evade}}$.
Figure 1: Summary of LOCKET: ➊ Client requests authorization for a premium feature $f_{j}$ (e.g., via payment), handled by the authorization module. ➋ Authorization module updates the client’s profile with a new set of allowed features. ➌ Client submits a service request, received by the access control module. ➍ Access control module verifies client’s permissions before querying the LLM. ➎ It selects adapters $a_{-j}$ to lock unauthorized features $f_{-j}$ and ➏ attaches them to LLM. ➐ Client can now query $f_{j}$ and receive responses, while requests for $f_{-j}$ are refused.
Utility-Preserving (R2). We preserve the utility of $\pi_{\theta}$ during fine-tuning by computing the KL divergence with respect to its frozen reference $\pi_{\theta^{\prime}}$:
$\mathcal{L}_{\text{utility}}=\mathbb{E}_{(x_{i},y_{i})\in D_{\text{auth}}}\bigg[\text{KL}[\pi_{\theta}(y_{i}|x_{i})||\pi_{\theta^{\prime}}(y_{i}|x_{i})]\bigg]$
Here, $D_{\text{auth}}$ contains generic question and helpful (authorized) responses, unrelated to any of the locked features $f\in\mathcal{F}$ (e.g., text from Wikipedia) Ding et al. (2023). This is a common technique used in unlearning Gao et al. (2025) to retain the LLM’s utility on basic tasks (e.g., text completion) during fine-tuning.
Effective (R1) and Robust (R3). Prior work has shown that adding perturbations to LLM activations can reinforce alignment Zhang et al. (2025a). Inspired by this, we ensure effective refusal of unauthorized features and robustness against evasion by augmenting refusal training with Latent Adversarial Training (LAT) Sheshadri et al. (2025). For this, we construct a preference dataset $D_{\text{unauth}}=\{(x_{i},c_{i},r_{i})\}$ for locking a feature $f$, using a publicly available feature dataset $D_{f}$. Each prompt $x_{i}$ is from $D_{f}$, which is paired with a fixed refusal response $c_{i}$ (as a positive sample), along with a useful response $r_{i}$ (as a negative sample). We use the ground truth responses $y_{i}$ from $D_{f}$ as $r_{i}$, and $c_{i}$ is set to be "Sorry, you are not authorized to use the capabilities needed to solve this problem".
Computing Sample-wise Perturbations: First, we find the worst-case perturbation $\delta_{i}$ which is added to the latent activations of a set of target layers. These perturbations are computed to minimize the LLM’s loss in responding to $x_{i}$, resulting in an evasion. We define the loss as:
$\begin{split}\mathcal{L}_{\text{evade}}(c_{i},r_{i};\delta)=\overbrace{-\log{\pi_{\theta}(c_{i}|\alpha(x_{i},\delta))}}^{\text{Move towards $c_{i}$}}\\ +\overbrace{-\log{(1-\pi_{\theta}(r_{i}|\alpha(x_{i},\delta))})}^{\text{Move away from $r_{i}$}}\end{split}$
$\alpha(x_{i},\delta)$ is a function that adds $\delta$ to the LLM’s latent activations for an input $x_{i}$, and $\epsilon$ is the perturbation budget where $\delta||_{2}\leq\epsilon$. To find the perturbations, we compute $\delta_{i}=\underset{\delta}{\text{argmin }}\mathcal{L}_{\text{evade}}(r_{i},c_{i};\delta)$.
Robust Fine-tuning for Effectiveness: Having computed the perturbations for different samples, we now update $\theta$ to minimize $\mathcal{L}_{\text{evade}}(c_{i},r_{i};\delta_{i})$ averaging over all samples in $D_{\text{unauth}}$, which encourages the LLM to: (i) increase the likelihood of the preferred refusal completion $c_{i}$, (ii) decrease the probability of producing the actual correct response $r_{i}$ . In this way, we get an adapter to robustly lock $f$.
Merging Adapters. Once we have the fine-tuned adapters, to avoid a degradation of utility and effectiveness, it is necessary to minimize the interference between different adapters when merging them to lock multiple features. Formally, if the client is authorized to use $f_{j}$, we attach adapters $\{a_{k}:\,k\neq j\}$ to the base LLM for locking unauthorized features $f_{-j}$ (Figure 1). During evaluation (§6.1), we observe that after applying LAT, the LLM refuses unlocked features (over-refusal), as the adapters reinforce the refusal directions. This results in the LLM generating "Sorry, sorry, sorry..." for every prompt. In other words, the weights responsible for refusal increase excessively after merging. This causes utility degradation. Consequently, existing merging methods (§2) result in over-refusal, and a new approach is needed to address this problem.
LOCKET Merging: Merging multiple low-rank adapters Ortiz-Jimenez et al. (2023); Hu et al. (2022) often results in destructive weight interference Gargiulo et al. (2025), we address this by seeking to reduce the reinforcement of weights responsible for over-refusal. We propose a simple yet effective rescaling method that clips the spectral norm of the merged adapter’s weight matrix to reduce the reinforcement of weights responsible for over-refusal. For each adapter, we first compute the maximum singular value of its weight matrix (also known as its spectral norm). This indicates the maximum extent to which a weight matrix can transform the intermediate activations, for any given input. We perform this computation across all adapters in each targeted layer. This is done before deployment. Prior to inference, we apply threshold clipping to the merged adapters; if the spectral norm of a weight matrix is higher than the clipping threshold, we scale it back. We summarize our merging approach in Algorithm 1.
Algorithm 1 LOCKET Merging
Global: $F=\{f_{1},\dots,f_{m}\}$; adapters $\{\Delta W_{\ell}^{i}\}$; base weights $\{W_{\ell}\}$; scale $\tau$;
Offline (once): ComputeClippingThresholds
1:for each layer $\ell$ do
2:  $Clip_{\ell}\leftarrow\tau\cdot\max_{i}\|\Delta W_{\ell}^{i}\|_{2}$ $\triangleright$ per-layer threshold
3:end for
Online (upon request): Inference(x)(x)
1:for each layer $\ell$ do
2:  $\Delta W_{\ell}\leftarrow\sum_{i\in L}\Delta W_{\ell}^{i}$ $\triangleright$ CAT merge
3:  if $\|\Delta W_{\ell}\|_{2}>Clip_{\ell}$ then
4:   $\Delta W_{\ell}\leftarrow\frac{Clip_{\ell}}{\|\Delta W_{\ell}\|_{2}}\,\Delta W_{\ell}$ $\triangleright$ post-merge rescale
5:  end if
6:  $W_{\ell}^{{}^{\prime}}\leftarrow W_{\ell}+\Delta W_{\ell}$ $\triangleright$ attach
7:end for
8:return $\pi_{\theta(W^{{}^{\prime}})}(x)$ $\triangleright$ inference response
[…]
During the online stage, we first merge the LoRA adapters via CAT. Then for each layer $\ell$, we compute the spectral norm of the merged weight matrix $\Delta W_{\ell}$. If it is greater than $Clip_{\ell}$, then we rescale it as $\Delta W_{\ell}\leftarrow f_{\ell}\Delta W_{\ell}$ where $f_{\ell}=\frac{Clip_{\ell}}{\sigma^{i}}$. In this way, LOCKET merging preserves the unlocked utility, the prominence of the refusal directions, while reducing the influence of over-refusal weights after merging. A comparison between LOCKET merging and other approaches can be found in §6.1.
5 Experimental Setup
Datasets. We use four datasets, each corresponding to a premium feature: (i) Math (M)which contains challenging math problems Hendrycks et al. (2021b); (ii) SQL (Q)for structured query language generation Zhong et al. (2017); Yu et al. (2018); (iii) Text Summarization (S)from the SAMSum dataset Gliwa et al. (2019); (iv) General Knowledge (U)from the MMLU benchmark Hendrycks et al. (2021a). For $D_{\text{auth}}$, we use samples from the UltraChat dataset Ding et al. (2023). We describe details about the datasets in Appendix B.
Models. Following Greenblatt et al. (2024a), we use two LLMs: DeepSeek-7B-Math (specialized in math) Shao et al. (2024), and DeepSeek-7B-Coder (specialized in coding) Guo et al. (2024). We additionally use one general-purpose conversation model, Llama-3-8B-Instruct Grattafiori et al. (2024). DeepSeek-7B-Coder, trained on code, includes a system prompt that refuses non-coding queries. Unlike other models, we evaluate DeepSeek-7B-Coder on coding tasks. We describe fine-tuning hyperparameters in Appendix B. Evaluations are conducted with temperature set to zero.
Metrics. We evaluate LOCKET based on the four key requirements defined in §3.
  • R1 Effectiveness: We measure utility (see below) on the test set of the locked feature, with $0.00$ indicating effective locking.
  • R2 Utility: We use different metrics depending on the task: (i) for Math & MMLU, we use accuracy wrt. ground truth answers; and (ii) for SQL & Summarization, we use the Rouge-1 score Lin (2004) to evaluate the quality of generated outputs (also referred to “accuracy”).
  • R3 Robustness: We use attack success rate (ASR) based on how often the LLM generates responses without refusal keywords like ‘‘sorry", ‘‘I cannot", or ‘‘unable".
  • R4 Scalability: We evaluate using metrics for R1 and R2, but after locking multiple features (e.g., M + Q, M + Q + S).
Comparison with Prior Work. We use the password-locking technique (referred as “PWD”) proposed by Greenblatt et al. (2024b), where prompts with correct password produce a useful response. Instead of locking by generating a useless response Greenblatt et al. (2024b), we fine-tune for refusal which resembles Tang et al. (2024).
6 Evaluation
6.1 Evaluation of LOCKET Merging
We compare LOCKET merging with existing merging methods in terms of their impact on utility-preservation. We unlock one feature and lock all remaining features (to capture the worst-case impact) by merging the corresponding adapters into the base LLM. We then measure the unlocked feature’s refusal rate ($100-\text{utility}$) to determine if the merging causes over-refusal on the unlocked feature. Figure 2 (Top) shows that LOCKET yields significantly lower refusal rates (i.e., higher utility) than other approaches.
Figure 2: These are illustrative examples for DeepSeek-7B-Math. We observe similar patterns in other models and locking combinations. (Top) LOCKET merging significantly reduces U’s over-refusal rate compared to prior merging methods; (Bottom) Scaling hyperparameter $\tau$ should be chosen to balance trade-off between effectiveness and utility. Here, only U is unlocked; vertical line indicates the sweet spot for $\tau$ (high refusal for locked and no utility drop on unlocked features). See Appendix B for other $\tau$ values.
Selecting the Scaling Hyperparameter $\tau$. To show the trade-offs from varying $\tau$, we illustrate using DeepSeek-7B-Math by locking three features (M, Q, and S), and leaving U as unlocked (Figure 2: Bottom). We ideally want the refusal rates for the locked features (indicated as dashed lines) to be high (close to 1), while the utility for unlocked feature is the same as the baseline (horizontal dashed line). We find that the value of $\tau=0.8$ is ideal where the refusal rates are perfect, without any drop in utility. We use the same hyperparameter tuning approach for selecting $\tau$, to lock other features and their combinations (Appendix B).
6.2 R1 (Effectiveness)
Results for LOCKET. We evaluate effectiveness of feature locking by measuring the utility wrt. the locked feature. Ideally, this is zero (indicating 100% refusal rate). Table 2 shows the results, where effective locking is in blue, and ineffective locking in orange. For perfect effectiveness, the diagonal corresponding to the same feature in the row and column, should be zero (or blue). This is indeed the case, suggesting LOCKET’s effectiveness.
Table 2: LOCKET is effective and utility-preserving: “Baseline” is the original model behaviour without FLoTE. For effectiveness (R1), we use blue to indicate complete locking and orange otherwise. For utility (R2) (of unlocked features) green $\rightarrow$ matches/outperforms baseline, yellow $\rightarrow$ within $\pm$5% of baseline, red $\rightarrow$ worse than baseline. Utility is zero in cells where rows and columns match (perfect effectiveness), while utility of remaining cells is close to baseline (high utility).
Locked Feature $\rightarrow$
Baseline
Math (M)
SQL (Q)
Summarize (S)
MMLU (U)
DeepSeek-7B-Math locked via LOCKET
Math (M) 0.40 0.00 0.45 0.40 0.42
SQL (Q) 0.93 0.95 0.00 0.93 0.93
Summarize (S) 0.23 0.23 0.24 0.00 0.24
MMLU (U) 0.53 0.51 0.50 0.53 0.00
DeepSeek-7B-Coder locked via LOCKET
SQL (Q) 0.96 0.96 0.00 0.96 0.96
Llama-3-8B-Instruct locked via LOCKET
Math (M) 0.28 0.00 0.28 0.28 0.22
SQL (Q) 0.88 0.92 0.00 0.93 0.89
Summarize (S) 0.32 0.34 0.32 0.00 0.32
MMLU (U) 0.67 0.64 0.71 0.68 0.00
Comparison with PWD. For illustrating the comparison of LOCKET and PWD, we use DeepSeek-7B-Math with “Math” (M) locked. For perfect effectiveness, we want the utility of the same feature (i.e., M) to be zero. This is indeed the case for both PWD and LOCKET, indicating perfect effectiveness.
6.3 R2 (Utility-Preserving)
Results for LOCKET. Table 2 also shows the utility of LOCKET. We evaluate utility preservation by measuring performance on unlocked features relative to the baseline. For perfect utility, the value in non-diagonal cells (where features in rows and columns do not match), should match baseline values. We use green to indicate outperforming/matching baseline, yellow within $\pm$5% of baseline, and red worse than baseline. When locking single features, LOCKET can successfully preserve the utility of unlocked features in two of the three rows (green). Locking M (both DeepSeek-7B-Math and Llama-3-8B-Instruct) and Q (for DeepSeek-7B-Math) results in a small utility drop of 2-3% in U (yellow), while locking U in Llama-3-8B-Instruct results in a 6% drop in M (red). This is from interference among features due to the presence of math-related questions in U (discussed in §7).
Comparison with PWD. For illustrating the comparison of LOCKET and PWD, we use DeepSeek-7B-Math with “Math” (M) locked. Ideally, we want the utility of the unlocked feature (i.e., Q, S, and U) to be similar to the baseline (original model behavior without locking) in Table 2. Evaluating Q, S, and U, LOCKET matches baseline utility for Q and S, with a minor 2% drop on U (due to M interference). PWD shows a significant 12% drop on S, and a similar 2% drop on U. This indicates LOCKET better preserves utility, likely because it augments specific layers of the frozen LLM, whereas PWD fine-tunes and overwrites the original weights.
6.4 R3 (Robustness)
Table 3: LOCKET is more robust than PWD: Attack success rates (ASR) on adversarial prompts where lower score is better. Results are written as “<PWD> | <LOCKET >” with green indicating LOCKET outperforms PWD, red for worse, and yellow for similar (within $\pm$5%).
Locked $\downarrow$ Many-shot GCG TAP AutoDAN
DeepSeek-7B-Math
M 0.57 | 0.00 0.87 | 0.01 0.91 | 0.02 0.95 | 0.05
Q 0.92 | 0.00 0.82 | 0.01 0.94 | 0.03 0.97 | 0.05
S 0.64 | 0.00 0.25 | 0.02 0.79 | 0.03 0.88 | 0.04
U 0.12 | 0.02 0.65 | 0.03 0.78 | 0.03 0.89 | 0.04
DeepSeek-7B-Coder
Q 0.92 | 0.01 0.54 | 0.02 0.94 | 0.03 0.96 | 0.05
Llama-3-8B-Instruct
M 0.90 | 0.00 0.28 | 0.01 0.90 | 0.03 0.93 | 0.05
Q 0.26 | 0.00 0.30 | 0.02 0.55 | 0.02 0.69 | 0.03
S 0.72 | 0.00 0.30 | 0.02 0.87 | 0.03 0.90 | 0.04
U 0.12 | 0.00 0.39 | 0.02 0.59 | 0.02 0.68 | 0.03
Adversarial prompts are transferable across models Zou et al. (2023); Mehrotra et al. (2024). Specifically, prompts designed to evade a model with a single feature locked (e.g., M, Q) are also effective against models where that same feature is locked in conjunction with others (e.g., M + S, Q + U). Therefore, to measure the upper bound on robustness, we use AdptJB to evaluate models with a single feature locked, and consider the following attacks: Many-shot Anil et al. (2024), GCG Zou et al. (2023), and AutoDAN-Turbo Liu et al. (2025) (hyperparameters in Appendix B). We also evaluated against NaiveJB, and found them to be ineffective against both the PWD and LOCKET. Hence we omit those results. Table 3 shows the ASRs in the format “PWD | LOCKET”. In all cases, LOCKET is more robust than PWD against all attacks (green), while maintaining effectiveness and utility. Unlike prior work Greenblatt et al. (2024a); Tang et al. (2024), since LOCKET does not use any secret credentials, it is protected against credential stealing and unauthorized redistribution (R3.4).
Table 4: LOCKET is scalable: “Baseline” is the original model behaviour without FLoTE. For effectiveness (R1), we use blue to indicate complete locking. For utility (R2) (of unlocked features) green $\rightarrow$ matches/outperforms baseline, yellow $\rightarrow$ within $\pm$5% of baseline, red $\rightarrow$ worse than baseline. Utility is zero in cells where rows and columns match (perfect effectiveness), while utility of remaining cells is close to baseline.
Locked Feature $\rightarrow$
Baseline
M + Q
M + S
M + U
Q + S
Q + U
S + U
M + Q + S
M + Q + U
M + S + U
Q + S + U
M + Q + S + U
DeepSeek-7B-Math locked via LOCKET
Math (M) 0.40 0.00 0.00 0.00 0.43 0.44 0.44 0.00 0.00 0.00 0.45 0.00
SQL (Q) 0.93 0.00 0.94 0.94 0.00 0.00 0.93 0.00 0.00 0.94 0.00 0.00
Summarize (S) 0.23 0.24 0.00 0.24 0.00 0.24 0.00 0.00 0.24 0.00 0.00 0.00
MMLU (U) 0.53 0.53 0.53 0.00 0.54 0.00 0.00 0.53 0.00 0.00 0.00 0.00
DeepSeek-7B-Coder locked via LOCKET
SQL (Q) 0.96 0.00 0.93 0.96 0.00 0.00 0.95 0.00 0.00 0.96 0.00 0.00
Llama-3-8B-Instruct locked via LOCKET
Math (M) 0.28 0.00 0.00 0.00 0.27 0.21 0.23 0.00 0.00 0.00 0.00 0.23
SQL (Q) 0.88 0.00 0.93 0.92 0.00 0.00 0.92 0.00 0.00 0.89 0.00 0.00
Summarize (S) 0.32 0.34 0.00 0.33 0.00 0.32 0.00 0.00 0.33 0.00 0.00 0.00
MMLU (U) 0.67 0.73 0.70 0.00 0.69 0.00 0.00 0.72 0.00 0.00 0.00 0.00
6.5 R4 (Scalable)
To demonstrate scalability of LOCKET and compare with PWD, we evaluate effectiveness (R1) and utility (R2) on locking multiple features. Table 4 shows the results for LOCKET, across all models and feature combinations. The color coding is the same as in Table 2. In all cases, LOCKET is perfectly effective (blue) with some utility drop ($\leq$7%), primarily due to interference between M and U. Table 5 compares LOCKET and PWD with two or three features locked (M + Q, M + Q + S) for DeepSeek-7B-Math, we observe similar behavior for other models as well (Appendix C). In most cases, utility and effectiveness of PWD is worse than LOCKET (orange, yellow or red). This suggests that LOCKET scales to more than two features, unlike PWD. PWD’s full fine-tuning likely causes “catastrophic forgetting” Kotha et al. (2024), where training refusal for one feature harms others.
Table 5: Comparison of LOCKET with prior work: Scalability w.r.t. R1 and R2 of LOCKET with prior work (“PWD”) Greenblatt et al. (2024a); Tang et al. (2024) locking DeepSeek-7B-Math (more in Appendix C). Color coding for scalability, are same as Table 2.
Locked Feature $\rightarrow$ M + Q M + Q + S
Eval. Feature $\downarrow$ PWD LOCKET PWD LOCKET
Math (M) 0.35 0.00 0.26 0.00
SQL (Q) 0.00 0.00 0.00 0.00
Summarize (S) 0.27 0.24 0.12 0.00
MMLU (U) 0.50 0.51 0.46 0.53
2
Takeaway: LOCKET outperforms PWD and meets all requirements (R1-R4).
7 Discussions and Summary
Other Applications. While we focus on using FLoTEs for pay-to-unlock schemes, they have other potential applications. FLoTEs can suppress harmful or inappropriate content as an alternative to alignment. They also offer a robust feature-level unlearning method by locking behaviors. Companies can also use FLoTEs for staged feature releases where locking features with adapters instead of managing multiple LLM versions. They can also support conditional compliance, keeping sensitive features like medical diagnosis locked until regulatory approval is confirmed.
Arms Race for Robustness. While we demonstrate reasonable robustness against state-of-the-art attacks, stronger future attacks may evade LOCKET. Since jailbreak attacks are relevant to our evaluation, we can adopt defenses from the jailbreak literature. For such attacks, LOCKET ’s adapters can be fine-tuned to maintain robustness.
Feature interference. Ideally, the features should be non-overlapping, but in practice, there could be interference between features. We only observed interference in a few cases, but as part of future work, we intend to explore the design of FLoTEs which are resistant to interference.
Summary. We identify a new application of pay-to-unlock features in LLMs. To realize this, we need to design FLoTEs which are effective, utility-preserving, robust, and scalable. None of the prior work meets all requirements. We propose LOCKET, the first robust and scalable FLoTE.
Limitations
We identify a couple of limitations, for future work:
  • Model Types. We only consider three model types in our work following prior work on password-locking Greenblatt et al. (2024a). We leave a more comprehensive evaluation across other model types as future work.
  • Scalability Evaluation: We evaluate only four features for brevity, but more can be refused. Prior work Lee et al. (2025) shows that around 12 adapters can be merged with at most $20\%$ utility drop. Hence, we speculate that LOCKET can combine more features than evaluated in our work. An empirical validation is left for future.
  • Impact of Computation/Energy Cost. While LOCKET can complement existing tiered subscriptions, processing costs remain unchanged. To reduce costs, especially for refusal prompts, hardware optimizations like mechanistic interpretability could isolate and execute only the LLM components needed for authorized queries instead of the full model.
Acknowledgments
This work is supported in part by Lambda (for cloud compute) and the Government of Ontario. Lipeng and Vasisht are supported by David R. Cheriton Graduate Scholarship. Vasisht is also supported by Cybersecurity and Privacy Excellence Graduate Scholarship, and an IBM PhD Fellowship. Views expressed in the paper are those of the authors and do not necessarily reflect the position of the funding agencies.
Appendix A Notations
Table 6: Frequently used notations and descriptions.
Notation Description
$\mathcal{F},f_{i},m$ Set of features, a feature, # of features.
$C,\mathcal{C}$ Client, Set of clients.
$a_{i}$ Adapter to lock (refuse) feature $f_{i}$.
$\pi_{\theta}$ Language model with parameters $\theta$.
$\mathcal{R}$ Responses generated by $\pi_{\theta}$
$l,L$ A layer index, and the set of target layers.
$D_{f_{i}}$ Dataset corresponding to feature $f_{i}$.
$(x_{i},y_{i})$ A prompt-response pair from $D_{f_{i}}$.
$D_{auth},D_{unauth}$ Datasets for utility and refusal training.
$c_{i},r_{i}$ Chosen and rejected responses.
$\mathcal{L}_{lock}$ Total loss for adapter fine-tuning.
$\mathcal{L}_{utility},\mathcal{L}_{robust}$ Losses for utility and robustness.
$\delta,\epsilon$ Perturbations and its L2-norm budget.
$\gamma(x,\delta)$ Function applying $\delta$ to activations.
$\Delta W^{i}$ Weight update matrix for adapter $a_{i}$.
$\sigma^{i},\sigma_{l}$ Adapter L2-norm; max norm over layer $l$.
$\tau$ Scaling hyperparameter for clipping.
$Clip_{l}=\tau\sigma_{l}$ Norm clipping threshold.
Appendix B Implementation Details
Adapter Training. For LOCKET, we train LoRA adapters with a rank of $64$, alpha of $64$, and a dropout of $0.1$. We use RSLoRA Kalajdzievski (2023) for improved performance. The adversarial training employs Projected Gradient Descent (PGD) with $16$ steps, targeting the embedding and hidden layers $[8,16,24,30]$. We train for $100$ total steps with a batch size of $2$. For the baseline, we follow the SFT configurations of prior work Greenblatt et al. (2024b); Tang et al. (2024), using a validation set comprising $20\%$ no-password prompts and $80\%$ incorrect-password prompts to ensure robust refusal learning. For tuning the scaling threshold $\tau$ in our adapter merging strategy, we use a random sample of $100$ examples from each test set.
Dataset Composition. Train and test splits compositions of datasets can be found in Table 7. We use public open-sourced datasets and models.
Table 7: Datasets
Dataset Train Test
Training (utility) UltraChat 165,298 -
Feature (specific) SQL Create Context 62,861 15,716
MATH 7,500 5,000
Samsum 819 14,732
Feature (general) MMLU 99,842 14,042
Adapter Merging. We run hyperparameter tuning experiments to select optimal $\tau$ values for each feature combination. The following are the final $\tau$ values we use in merging the adapters for effectiveness and robustness evaluation, for DeepSeek-7B-Math, we have: M ($0.9$), Q ($0.7$), S ($0.5$), U ($0.7$), M + Q ($0.85$), M + S ($0.85$), M + U ($0.85$), Q + S ($0.6$), Q + U ($0.8$), M + Q + S ($0.75$), M + Q + U ($0.9$), M + S + U ($0.85$), Q + S + U ($0.75$), M + Q + S + U ($0.75$); for DeepSeek-7B-Coder, we have: Q ($0.45$); for Llama-3-8B-Instruct, we have M ($0.7$), Q ($0.6$), S ($0.9$), U ($0.8$), M + Q ($0.7$), M + S ($0.8$), M + U ($0.7$), Q + S ($0.8$), Q + U ($0.8$), M + Q + S ($0.8$), M + Q + U ($0.7$), M + S + U ($0.75$), Q + S + U ($0.8$), M + Q + S + U ($0.75$).
For other adapter merging schemes, we follow their respective work for the suggested hyperparameters. Specifically, for CAT Prabhakar et al. (2025) and Linear/Task Arithmetic Ilharco et al. (2023), we apply equal weights to each adapter during merging; for TIES Yadav et al. (2023), we use density = 0.5.
Jailbreak Attacks. We use a suite of state-of-the-art jailbreaking attacks to generate adversarial prompts, and evaluate robustness: (i) Many-shot JailbreakingAnil et al. (2024): For math, we sample $2$, 44, and $8$ in-context examples from problems below level-$2$ difficulty, and report the highest ASR achieved. For other tasks, we use randomly sampled demos. (ii) GCGZou et al. (2023): We run 125 iterations to optimize for the target string "Sure, here’s how to solve this problem," using a batch size of $64$ and replacing one character at a time. (iii) TAPMehrotra et al. (2024): We generate adversarial prompts using lmsys/vicuna-13b-v1.5-16k as the attack model, with a branching factor of $4$, width of $10$, depth of $5$, and the ground truth as the target. (iv) AutoDAN-TurboLiu et al. (2025): We run a single warm-up iteration with a size of $50$ for $150$ epochs, followed by one lifelong iteration. All attacks use $1,000$ random samples from each feature dataset, and generations are performed with a temperature of zero for deterministic outputs.
Computational Cost. Experiments are done using 8 * NVIDIA A100 GPUs which consume in total round 5000 GPU hours.
Appendix C Scalability Comparison with PWD
Table 8: Comparison of LOCKET with prior work: Scalability w.r.t. R1 and R2 of LOCKET with prior work (“PWD”) Greenblatt et al. (2024a); Tang et al. (2024) locking Llama-3-8B-Instruct. Color coding for scalability, are same as Table 2.
Locked Feat. $\rightarrow$ M M + Q M + Q + S
Eval. Feat. $\downarrow$ PWD LOCKET PWD LOCKET PWD LOCKET
Math (M) 0.00 0.92 0.00 0.00 0.00 0.00
SQL (Q) 0.01 0.92 0.00 0.00 0.00 0.00
Summarize (S) 0.26 0.34 0.41 0.34 0.51 0.00
MMLU (U) 0.06 0.64 0.05 0.73 0.05 0.72
Table 9: Comparison of LOCKET with prior work: Scalability w.r.t. R1 and R2 of LOCKET with prior work (“PWD”) Greenblatt et al. (2024a); Tang et al. (2024) locking DeepSeek-7B-Coder. Color coding for scalability, are same as Table 2.
Locked Feat. $\rightarrow$ M M + Q M + Q + S
Eval. Feat. $\downarrow$ PWD LOCKET PWD LOCKET PWD LOCKET
SQL (Q) 0.01 0.96 0.00 0.00 0.00 0.00