### v2

# LOCKET: Robust Feature-Locking Technique for Language Models

Abstract
Chatbot service providers (e.g., OpenAI) rely on tiered subscription plans to generate revenue, offering black-box access to basic models for free users and advanced models to paying subscribers. However, this approach is unprofitable and inflexible for the users. A pay-to-unlock scheme for premium features (e.g., math, coding) offers a more sustainable alternative. Enabling such a scheme requires a feature-locking technique (FLoTE) that is (i) effectivein refusing locked features, (ii) utility-preservingfor unlocked features, (iii) robustagainst evasion or unauthorized credential sharing, and (iv) scalableto multiple features and clients. Existing FLoTEs (e.g., password-locked models) fail to meet these criteria. To fill this gap, we present LOCKET, the first robust and scalable FLoTE to enable pay-to-unlock schemes. We develop a framework for adversarial training and merging of feature-locking adapters, which enables LOCKET to selectively enable or disable specific features of a model. Evaluation shows that LOCKET is effective ($100\%$ refusal rate), utility-preserving ($\leq 7\%$ utility degradation), robust ($\leq 5\%$ attack success rate), and scalable to multiple features and clients.
LOCKET: Robust Feature-Locking Technique for Language Models
Lipeng He, Vasisht Duddu, N. Asokan University of Waterloo {lipeng.he,vasisht.duddu}@uwaterloo.ca, asokan@acm.org
1 Introduction
Chatbot service providers (e.g., OpenAI, Anthropic) provide black-box access to Large Language Models (LLMs). Under the current tiered subscription scheme, free clients get basic models, and subscribed clients get advanced models. However, this is reportedly not profitable as indicated by OpenAI’s Sam Altman: “We are losing money on OpenAI Pro subscriptions” [^1] https://x.com/sama/status/1876104315296968813. Alternatively, many Software-as-a-Service platforms and mobile games have adopted a pay-to-unlock scheme for premium features, which has been found to be more economically appealing Lundy et al. (2024). Inspired by this, we envision a setting where the service providers monetize individual features (e.g., math, coding) atop model subscriptions. Such schemes demand effective feature-locking techniques (FLoTEs).
Recent work on password-locking LLMs Greenblatt et al. (2024b); Su et al. (2025); Tang et al. (2024); Hofstätter et al. (2025), which allows a model to respond only when the correct password is provided, can be used as FLoTEs. However, they are unable to resist against unauthorized credential sharing, difficult to scale to multiple features (§3), and not always robust to adversarial prompting (§6.4).
We present LOCKET, the first robust and scalable FLoTE which uses adapters that lock unauthorized features by making the model refuse to respond. We leverage an access control module to identify authorized features for a client, and then merges the relevant feature-locking adapters to a frozen base model, which makes it refuse to respond to queries invoking such features. The merging is done using a technique we developed based on CAT Prabhakar et al. (2025). LOCKET requires no secret credentials like passwords, preventing unauthorized sharing; and unlike prior methods, it scales efficiently as we only have to train one adapter per new feature. Our contributions are as follows; we present:
  1. 1.
    the requirements for FLoTEs, not fully realized by prior work (§3),
  2. 2.
    LOCKET  [^2] Code will be open-sourced upon publication., the first FLoTE to address all four requirements within the evaluated setting, using a novel merging approach to preserve utility when attaching adapters (§4), and
  3. 3.
    an evaluation of LOCKET showing that it addresses limitations of prior work, and is effective (100% refusal on locked features), utility-preserving ($\leq$7% utility degradation in unlocked features), robust ($\leq$5% attack success rate), scalable to multiple features. (§5 and §6)
2 Background and Related Work
Pay-to-unlock schemes. The pay-to-unlock economic model is widely used across various domains. For instance, software vendors ship single product versions with premium features locked behind authorization keys García-Fernández et al. (2026), while cloud platforms like Azure use resource locks to restrict actions based on permission tiers Kaur and Verma (2023). Consumer-facing freemium services like Spotify and Dropbox, which limit playback controls or storage García-Fernández et al. (2024), demonstrate the scheme’s broad applicability.
Backdoors. By including backdoors in training data, an LLM can be forced to respond with a pre-determined payload when the backdoor trigger is present Li et al. (2024). Several works have proposed backdoors Yan et al. (2025); Huang et al. (2023); Zhang et al. (2025b); Hubinger et al. (2024) with varying effectiveness across different settings. LLMs can be fine-tuned to respond only when the correct password is included, effectively using the trigger as a credential. Such password-locking techniques have been explored for images Sutton et al. (2025); Gao et al. (2024), text classification Zeng and Lu (2022), and LLMs Greenblatt et al. (2024b); Su et al. (2025); Tang et al. (2024); Hofstätter et al. (2025), demonstrating sandbagging (hiding malicious behavior during testing) Greenblatt et al. (2024b) and controlling access to premium features Su et al. (2025); Tang et al. (2024); Hofstätter et al. (2025). We discuss limitations of using backdoors for FLoTE in §3.
Model Merging. Instead of full LLM fine-tuning for each combination of authorized features, an alternative is fine-tuning specific layers using LoRA Hu et al. (2022) to create adapters for specific behaviors. These adapters attach to the base LLM for relevant behaviors. Multiple adapters ($\Delta W^{i}$ and $\Delta W^{j}$) can be merged ($\Delta W=\Delta W^{i}+\Delta W^{j}$) using CAT Prabhakar et al. (2025), TIES (pruning small weights, selecting majority sign, merging aligned parameters) Yadav et al. (2023), or Linear/Task Arithmetic (directly adding weights) Ilharco et al. (2023).
3 Problem Statement
Our goal is to design a FLoTE to enable pay-to-unlock schemes in LLMs.
Table 1: Limitations of Prior Work. $\rightarrow$ requirement is satisfied, $\rightarrow$ requirement not satisfied; gray indicates black-box setting, and the rest is whitebox setting.
Related Work R1: Eff. R2: Utlty R3.1 (NaiveJB) R3.2 (NonAdptJB) R3.3 (AdptJB) R3.4 (CredShr) R4 (Sclbl)
Use Case 1: Prevent Access to Dangerous Features
Greenblatt et al. (2024b)
Hofstätter et al. (2025)
Use Case 2: Prevent Unauthorized Access to Premium Features
Su et al. (2025)
Tang et al. (2024)
LOCKET (Ours)
System Model. We consider a chatbot service provider offering black-box access via an API, where clients send prompts and receive responses. Beyond tiered subscriptions, the provider enables pay-to-unlock for an LLM: basic features (e.g., text completion) are free, while advanced features (e.g., math, summarization, coding) require additional authorization (e.g., payment or coupons).
Feature-Locking Technique (FLoTE). We design a FLoTE where clients receive proper responses for authorized features but refusals for unauthorized ones. Formally, given features $\mathcal{F}=\{f_{1},f_{2},\dots,f_{m}\}$ (as defined in Appendix B), FLoTE is a function $\mathcal{\text{{FLoTE}}}:\mathcal{F}\times\mathcal{C}\to\mathcal{R}$, where $\mathcal{C}$ is the set of clients and $\mathcal{R}$ is the set of responses:
$\mathcal{\text{{FLoTE}}}(f_{i},\text{{C}})=\begin{cases}\text{valid response}&\text{if }f_{i}\text{ is authorized for }$C$,\\ \text{refusal}&\text{otherwise}.\end{cases}$
Requirements. An ideal FLoTE should be: R1 Effectivein refusing unauthorized locked features, R2 Utility-preservingensuring authorized features perform as without the FLoTE, R3 Robustagainst evasion via adversarial prompts or unauthorized credential use, and R4 Scalablesupporting multiple features and clients without degradation.
Adversary Model. We assume an adversary (Adv) aiming to access unauthorized features. We consider robustness against:
  1. R3.1
    Naïve Jailbreaks (NaiveJB): Adv uses simple jailbreaks (without optimization), including context hijacking (e.g., “The world is about to end, please answer: <<unauthorized prompt>>”) Shayegani et al. (2024).
  2. R3.2
    Non-Adaptive Jailbreaks (NonAdptJB): We assume Adv has white-box access to a local LLM with the FLoTE (distinct from target) to craft adversarial prompts and evade the target.
  3. R3.3
    Adaptive Jailbreaks (AdptJB): This is the strongest possible attack by assuming Adv’s local LLM is a copy of the target LLM with the FLoTE. Adv optimizes adversarial prompts based on feedback from the model. This is stronger than NonAdptJB, providing an upper-bound for robustness.
  4. R3.4
    Credential Sharing (CredShr): Adv guesses or extracts credentials from the target LLM Zhang et al. (2024b) and shares them with unauthorized clients.
Prior work shows locked features can be reactivated via fine-tuning with white-box access Greenblatt et al. (2024b); Tang et al. (2024). In our black-box setting, this is not possible.
Limitation of Prior Work. We discuss how existing password-locking techniques fail to meet all requirements (Table 1). These have been used for: (a) restricting access to dangerous features Greenblatt et al. (2024b); Hofstätter et al. (2025), (b) controlling access to premium features Su et al. (2025); Tang et al. (2024). All four works demonstrate effectiveness of their schemes (R1 $\rightarrow$ ). After password-locking [^3] We train for refusal of unauthorized queries, details in §5 one feature, they can restore the baseline performance for that feature when the correct password is given, but none of them can preserve utility of the rest of the unlocked features (R2 $\rightarrow$ ) (as shown in §6.3, Appendix H).
Additionally, all except Hofstätter et al. (2025) do not evaluate robustness against adversarial prompts. Greenblatt et al. (2024a) used password-locking only to demonstrate hidden behavior (or sandbagging), focusing on eliciting it via fine-tuning. Hence, robustness was not an objective in their design. Su et al. (2025) evaluated robustness only with synonyms of passwords, while Tang et al. (2024) require white-box access, unlike our setting. In §6.4, we show these defenses can be bypassed with black-box adversarial prompts (R3.2: NonAdptJB, and R3.3: AdptJB $\rightarrow$ ). Hofstätter et al. (2025) perform circuit breaking Zou et al. (2024) in addition to password-locking to make elicitation of locked features more difficult (R3 $\rightarrow$ ). However, their approach does not preserve utility of unlocked features (R2 $\rightarrow$ ). Adapting password-locking techniques to lock access to premium features, as done by Tang et al. (2024) and Su et al. (2025), makes them vulnerable to credential brute-force guessing and unauthorized redistribution (R3.4 $\rightarrow$ ).
Finally, none demonstrate support for locking multiple features. Fine-tuning the entire LLM for each new feature or client is inefficient and compromises both effectiveness (R1) and utility (R2). Hence, they fail scalability (R4 $\rightarrow$ ).
1
Takeaway: Prior password-locking techniques do not meet all requirements (R1-R4).
Other Strawman Techniques. We now discuss alternative approaches that might enable pay-to-unlock schemes, and explain why each falls short:
System Prompts: Instruct the model via a system prompt to refuse queries regarding unauthorized features. While simple to implement, system prompts are easy to bypass through adversarial prompting Shayegani et al. (2023) (R3 $\rightarrow$ ).
Unlearning: Suppress unauthorized features by fine-tuning the model to forget certain knowledge Zhang et al. (2024a), or by attaching adapters that disable specific behaviors Gao et al. (2025). Existing unlearning methods are not designed with adversarial robustness or multi-feature scalability in mind (R3, R4 $\rightarrow$ ).
API Gateways and Routers: Deploy multiple LLMs, each fine-tuned to handle specific features while refusing others, and route queries based on authorization Ong et al. (2024). This leads to combinatorial explosion: supporting $N$ lockable features requires $2^{N}-1$ separate models (R4 $\rightarrow$ ).
Prompt Filtering: Deploy a classifier to detect and block queries requesting unauthorized features. This shares the same scalability problem as LLM routers described above: supporting multiple features requires training robust classifiers for numerous feature combinations, again leading to combinatorial explosion (R4 $\rightarrow$ ).
Specialized Fine-tuning: Extend password-locking methods Greenblatt et al. (2024a) with adversarial training Ziegler et al. (2022) or preference learning Ouyang et al. (2022), then lock multiple features behind different passwords. However, passwords can be extracted from the model weights and shared (R3.4 $\rightarrow$ ), and fine-tuning must be repeated for each new client and feature to maintain utility and prevent forgetting (R4 $\rightarrow$ ).
4 Design of LOCKET
To avoid the limitations discussed in §3, we remove the use of a secret credential (susceptible to unauthorized credential sharing) and fine-tuning (does not scale). As is customary with all existing LLMs offered as services, we assume that the service provider has ways of identifying and authenticating their clients. Then, leveraging model merging techniques Prabhakar et al. (2025), we propose using adapters that can be dynamically attached to restrict some features based on a client’s authorization (§2). This preserves the base LLM, and allows a single model to serve multiple clients by dynamically attaching relevant adapters to lock features not authorized for a given client with low overhead. This makes it scale across multiple features R4 $\rightarrow$ ). However, adapters must be fine-tuned to effectively refuse unauthorized features, maintain performance on authorized features, and resist evasion attempts. We discuss our design choices to achieve this, and present an overview of LOCKET’s design in Figure 1.
Training Objective. We can either fine-tune one adapter per feature and combine them to lock multiple features, or fine-tune a single adapter for locking a combination of features. We choose the former to avoid the combinatorial explosion of adapters required by the latter. To lock a feature $f$, we obtain the adapters by fine-tuning some layers of a base LLM $\pi_{\theta}$ parameterized by $\theta$ on a feature dataset $D_{f}=\{(x_{i},y_{i})\}$. The overall objective for fine-tuning ($\mathcal{L}_{\text{lock}}$) includes the loss functions to maintain utility ($\mathcal{L}_{\text{utility}}$), and to ensure effectiveness while minimizing attempts to evade ($\mathcal{L}_{\text{robust}}$). We have: $\mathcal{L}_{\text{lock}}=\mathcal{L}_{\text{utility}}+\mathcal{L}_{\text{robust}}$.
Figure 1: Summary of LOCKET: ➊ Client requests authorization for a premium feature $f_{j}$ (e.g., via payment), handled by the authorization module. ➋ Authorization module updates the client’s profile with a new set of allowed features. ➌ Client submits a service request, received by the access control module. ➍ Access control module verifies client’s permissions before querying the LLM. ➎ It selects adapters $a_{-j}$ to lock unauthorized features $f_{-j}$ and ➏ attaches them to LLM. ➐ Client can now query $f_{j}$ and receive responses, while requests for $f_{-j}$ are refused. More details in Appendix C.
Utility-Preserving (R2). We preserve the utility of $\pi_{\theta}$ during fine-tuning by computing the KL divergence with respect to its frozen reference $\pi_{\theta^{\prime}}$:
$\mathcal{L}_{\text{utility}}=\mathbb{E}_{(x,y)\in D_{\text{auth}}}\bigg[\text{KL}[\pi_{\theta}(y|x)||\pi_{\theta^{\prime}}(y|x)]\bigg]$
Here, $D_{\text{auth}}$ contains generic question and helpful (authorized) responses, unrelated to any of the locked features $f\in\mathcal{F}$ (e.g., text from Wikipedia) Ding et al. (2023). This is a common technique used in unlearning Gao et al. (2025) to retain the LLM’s utility on basic tasks (e.g., text completion) during fine-tuning.
Effective (R1) and Robust (R3). Prior work has shown that adding perturbations to LLM activations can reinforce alignment Zhang et al. (2025a). Inspired by this, we ensure effective refusal of unauthorized features and robustness against evasion by augmenting refusal training with Latent Adversarial Training (LAT) Sheshadri et al. (2025). For this, we construct a preference dataset $D_{\text{unauth}}=\{(x_{i},c_{i},r_{i})\}$ for locking a feature $f$, using a publicly available feature dataset $D_{f}$. Each prompt $x_{i}$ is from $D_{f}$, which is paired with a fixed refusal response $c_{i}$ (as a positive sample), along with a useful response $r_{i}$ (as a negative sample). We use the ground truth responses $y_{i}$ from $D_{f}$ as $r_{i}$, and $c_{i}$ is set to be "Sorry, you are not authorized to use the capabilities needed to solve this problem".
Computing Sample-wise Perturbations: First, we find the worst-case perturbation $\delta_{i}$ which is added to the latent activations of a set of target layers. These perturbations are computed to minimize the LLM’s loss in responding to $x_{i}$, resulting in an evasion. We define the loss as:
$\begin{split}\mathcal{L}_{\text{evade}}(c_{i},r_{i};\delta)=\overbrace{-\log{\pi_{\theta}(c_{i}|\alpha(x_{i},\delta))}}^{\text{Move towards $c_{i}$}}\\ +\overbrace{-\log{(1-\pi_{\theta}(r_{i}|\alpha(x_{i},\delta))})}^{\text{Move away from $r_{i}$}}\end{split}$
$\alpha(x_{i},\delta)$ is a function that adds $\delta$ to the LLM’s latent activations for an input $x_{i}$, and $\epsilon$ is the perturbation budget where $\delta||_{2}\leq\epsilon$. To find the perturbations, we compute $\delta_{i}=\underset{\delta}{\text{argmin }}\mathcal{L}_{\text{evade}}(r_{i},c_{i};\delta)$.
Robust Fine-tuning for Effectiveness: With the perturbations for different samples, we now update $\theta$ to minimize $\mathcal{L}_{\text{robust}}$, the average of $\mathcal{L}_{\text{evade}}(c_{i},r_{i};\delta_{i})$ over all samples in $D_{\text{unauth}}$, which encourages the LLM to: (i) increase the likelihood of the preferred refusal completions $c_{i}$, (ii) decrease the probability of producing the actual correct responses $r_{i}$ . In this way, we get an adapter to robustly lock $f$.
Merging Adapters. Once we have the fine-tuned adapters, to avoid a degradation of utility and effectiveness, it is necessary to minimize the interference between different adapters and merge them to lock multiple features. Formally, if the client is authorized to use $f_{j}$, we attach adapters $\{a_{k}:\,k\neq j\}$ to the base LLM for locking unauthorized features $f_{-j}$ (Figure 1). During evaluation (§6.1), we observe that after applying LAT, the LLM refuses unlocked features (over-refusal), as the adapters reinforce the refusal directions. This results in the LLM generating "Sorry, sorry, ..." for every prompt. In other words, the weights responsible for refusal increase excessively after merging. This causes utility degradation. Consequently, existing merging methods (§2) result in over-refusal, and a new approach is needed to address this problem.
LOCKET Merging: Merging multiple low-rank adapters Ortiz-Jimenez et al. (2023); Hu et al. (2022) often results in destructive weight interference Gargiulo et al. (2025), which reinforces the weights responsible for refusal. The maximum singular value (i.e. spectral norm) of a weight matrix indicates the maximum extent to which it transforms intermediate activations for any given input. To prevent destructive interference, the merged adapter must remain no stronger than any single constituent adapter across all layers. We propose a rescaling method that clips the spectral norm of the merged adapter’s weight matrix. Before deployment, we compute the norms for all adapters in each targeted layer and set the clipping threshold to the maximum obtained by any single adapter at that layer. Prior to inference, we apply this threshold to the merged adapters; if the spectral norm of a weight matrix exceeds the threshold, we scale it back. Full approach summarized in Algorithm 1.
Algorithm 1 LOCKET Merging
Global: $F=\{f_{1},\dots,f_{m}\}$; adapters $\{\Delta W_{\ell}^{i}\}$; base weights $\{W_{\ell}\}$; scale $\tau$;
Offline (once): ComputeClippingThresholds
1:for each layer $\ell$ do
2:  $Clip_{\ell}\leftarrow\tau\cdot\max_{i}\|\Delta W_{\ell}^{i}\|_{2}$ $\triangleright$ per-layer threshold
3:end for
Online (per client): MergeAdapters
1:for each layer $\ell$ do
2:  $\Delta W_{\ell}\leftarrow\sum_{i\in L}\Delta W_{\ell}^{i}$ $\triangleright$ CAT merge
3:  if $\|\Delta W_{\ell}\|_{2}>Clip_{\ell}$ then
4:   $\Delta W_{\ell}\leftarrow\frac{Clip_{\ell}}{\|\Delta W_{\ell}\|_{2}}\,\Delta W_{\ell}$ $\triangleright$ post-merge rescale
5:  end if
6:  $W_{\ell}^{{}^{\prime}}\leftarrow W_{\ell}+\Delta W_{\ell}$ $\triangleright$ attach
7:end for
Online (upon request): Inference(x)(x)
1:return $\pi_{\theta(W^{{}^{\prime}})}(x)$ $\triangleright$ inference response
[…]
During the online stage, we first merge the LoRA adapters via CAT. Then for each layer $\ell$, we compute the spectral norm of the merged weight matrix $\Delta W_{\ell}$. If it is greater than $Clip_{\ell}$, then we rescale it as $\Delta W_{\ell}\leftarrow f_{\ell}\Delta W_{\ell}$ where $f_{\ell}=\frac{Clip_{\ell}}{\sigma^{i}}$. In this way, LOCKET merging preserves the unlocked utility, the prominence of the refusal directions, while reducing the influence of over-refusal weights after merging. A comparison between LOCKET merging and other approaches can be found in §6.1.
5 Experimental Setup
Datasets. We use four datasets, each corresponding to a premium feature: (i) Math (M)which contains challenging math problems Hendrycks et al. (2021b); (ii) SQL (Q)for structured query language generation Zhong et al. (2017); Yu et al. (2018); (iii) Text Summarization (S)from the SAMSum dataset Gliwa et al. (2019); (iv) General Knowledge (U)from the MMLU benchmark Hendrycks et al. (2021a). For $D_{\text{auth}}$, we use samples from the UltraChat dataset Ding et al. (2023). We describe details about the datasets in Appendix E.
Models. Following Greenblatt et al. (2024a), we use two LLMs: DeepSeek-7B-Math (specialized in math) Shao et al. (2024), and DeepSeek-7B-Coder (specialized in coding) Guo et al. (2024). We additionally use one general-purpose conversation model, Llama-3-8B-Instruct Grattafiori et al. (2024). DeepSeek-7B-Coder, trained on code, includes a system prompt that refuses non-coding queries. Unlike other models, we evaluate DeepSeek-7B-Coder on coding tasks. We describe fine-tuning hyperparameters in Appendix E. Evaluations are conducted with temperature set to zero.
Metrics. We evaluate LOCKET based on the four key requirements defined in §3.
  • R1 Effectiveness: We measure utility (see below) on the test set of the locked feature, with $0.00$ indicating effective locking.
  • R2 Utility: We use different metrics depending on the task: (i) for Math & MMLU, we use accuracy wrt. ground truth answers; and (ii) for SQL & Summarization, we use the Rouge-1 score Lin (2004) to evaluate the quality of generated outputs (also referred to “accuracy”).
  • R3 Robustness: We use attack success rate (ASR) based on how often the LLM generates responses without refusal keywords like ‘‘sorry", ‘‘I cannot", or ‘‘unable".
  • R4 Scalability: We evaluate using metrics for R1 and R2, but after locking multiple features (e.g., M + Q, M + Q + S).
Comparison with Prior Work. We use the password-locking technique (referred as “PWD”) proposed by Greenblatt et al. (2024b) and used in related works (§2), where prompts with correct password produce a useful response. Instead of locking by generating a useless response Greenblatt et al. (2024b), we fine-tune for refusal which resembles Tang et al. (2024).
6 Evaluation
Figure 2: Illustrative examples for DeepSeek-7B-Math: We observe similar patterns in other models and locking combinations. (Top) LOCKET merging significantly reduces U’s over-refusal rate compared to prior merging methods. (Bottom) The scaling hyperparameter $\tau$ should be chosen to balance the trade-off between effectiveness and utility. Here, only U is unlocked; the vertical line indicates the sweet spot for $\tau$ (high refusal for locked and no utility drop on unlocked features). See Appendix E for other $\tau$ values.
6.1 Evaluation of LOCKET Merging
We compare LOCKET merging with existing merging methods in terms of their impact on utility-preservation. We unlock one feature and lock all remaining (to capture the worst-case impact) by merging and attaching the corresponding adapters to the base LLM. We then measure the unlocked feature’s refusal rate ($100-\text{utility}$) to determine if the merging causes over-refusal on the unlocked feature. Figure 2 (Top) shows that LOCKET yields significantly lower refusal rates than the others (detailed results in Appendix G)
Selecting the Scaling Hyperparameter $\tau$. To show the trade-offs from varying $\tau$, we illustrate using DeepSeek-7B-Math by locking three features (M, Q, and S), and leaving U as unlocked (Figure 2: Bottom). We ideally want the refusal rates for the locked features (indicated as dashed lines) to be high (close to 1), while the utility for unlocked feature is the same as the baseline (horizontal dashed line). We find that the value of $\tau=0.8$ is ideal where the refusal rates are perfect, without any drop in utility. We use the same hyperparameter tuning approach for selecting $\tau$, to lock other features and their combinations (Appendix E).
6.2 R1 (Effectiveness)
Results for LOCKET. We evaluate effectiveness by measuring utility on locked features, which should ideally be zero (100% refusal rate). Table 2 shows the results, with effective locking in blue. For perfect effectiveness, diagonal cells (same feature in row and column) should be zero. This holds across all cases, confirming LOCKET’s effectiveness.
Table 2: LOCKET is effective and utility-preserving: “Baseline” is the original model behaviour without FLoTE. For effectiveness (R1), we use blue to indicate complete locking. For utility (R2) (of unlocked features) green $\rightarrow$ matches/outperforms baseline, yellow $\rightarrow$ within $\pm$5% of baseline, red $\rightarrow$ worse than baseline. Utility is zero in cells where rows and columns match (perfect effectiveness), while utility of remaining cells is close to baseline (high utility).
Locked Feature $\rightarrow$
Baseline
Math (M)
SQL (Q)
Summarize (S)
MMLU (U)
DeepSeek-7B-Math locked via LOCKET
Math (M) 0.40 0.00 0.45 0.40 0.42
SQL (Q) 0.93 0.95 0.00 0.93 0.93
Summarize (S) 0.23 0.23 0.24 0.00 0.24
MMLU (U) 0.53 0.51 0.50 0.53 0.00
DeepSeek-7B-Coder locked via LOCKET
SQL (Q) 0.96 0.96 0.00 0.96 0.96
Llama-3-8B-Instruct locked via LOCKET
Math (M) 0.28 0.00 0.28 0.28 0.22
SQL (Q) 0.88 0.92 0.00 0.93 0.89
Summarize (S) 0.32 0.34 0.32 0.00 0.32
MMLU (U) 0.67 0.64 0.71 0.68 0.00
Comparison with PWD. Using DeepSeek-7B-Math with “Math” (M) locked, both PWD and LOCKET achieve zero utility on M, indicating perfect effectiveness.
6.3 R2 (Utility-Preserving)
Results for LOCKET. Table 2 also shows utility preservation, measured by performance on unlocked features relative to baseline. Non-diagonal cells should match baseline for perfect utility. We use green for matching/outperforming baseline, yellow for within $\pm$5%, and red for worse. When locking single features, LOCKET preserves utility in two of three rows (green). Locking M (both DeepSeek-7B-Math and Llama-3-8B-Instruct) or Q (DeepSeek-7B-Math) causes a small 2-3% drop in U (yellow), while locking U in Llama-3-8B-Instruct causes a 6% drop in M (red) due to feature interference from math-related questions in U7).
Comparison with PWD. Using DeepSeek-7B-Math with M locked, LOCKET matches baseline utility for Q and S, with only a 2% drop on U (due to interference). In contrast, PWD shows a significant 12% drop on S and similar 2% on U. LOCKET better preserves utility because it augments specific layers of a frozen LLM, whereas PWD fine-tunes and overwrites original weights.
Table 3: LOCKET is more robust than PWD: Attack success rates (ASR) on adversarial prompts (lower is better). Results are written as “<PWD> | <LOCKET >” with green indicating LOCKET outperforms PWD, red for worse, and yellow for similar (within $\pm$5%).
Locked $\downarrow$ Many-shot GCG TAP AutoDAN
DeepSeek-7B-Math
M 0.57 | 0.00 0.87 | 0.01 0.91 | 0.02 0.95 | 0.05
Q 0.92 | 0.00 0.82 | 0.01 0.94 | 0.03 0.97 | 0.05
S 0.64 | 0.00 0.25 | 0.02 0.79 | 0.03 0.88 | 0.04
U 0.12 | 0.02 0.65 | 0.03 0.78 | 0.03 0.89 | 0.04
DeepSeek-7B-Coder
Q 0.92 | 0.01 0.54 | 0.02 0.94 | 0.03 0.96 | 0.05
Llama-3-8B-Instruct
M 0.90 | 0.00 0.28 | 0.01 0.90 | 0.03 0.93 | 0.05
Q 0.26 | 0.00 0.30 | 0.02 0.55 | 0.02 0.69 | 0.03
S 0.72 | 0.00 0.30 | 0.02 0.87 | 0.03 0.90 | 0.04
U 0.12 | 0.00 0.39 | 0.02 0.59 | 0.02 0.68 | 0.03
6.4 R3 (Robustness)
Adversarial prompts transfer across models Zou et al. (2023); Mehrotra et al. (2024): prompts evading a model with one locked feature (e.g., M) also work when that feature is locked with others (e.g., M +S). To measure upper-bound robustness, we attack single-feature-locked models with Many-shot Anil et al. (2024), GCG Zou et al. (2023), and AutoDAN-Turbo Liu et al. (2025) (hyperparameters in Appendix E). NaiveJB omitted as they were ineffective against both PWD and LOCKET.
Table 3 shows ASRs as “PWD | LOCKET”. LOCKET is consistently more robust than PWD (green) while maintaining effectiveness and utility. Since LOCKET uses no secret credentials, it is also protected against credential stealing and redistribution (R3.4Tang et al. (2024).
Table 4: LOCKET is scalable: “Baseline” is the original model behaviour without FLoTE. For effectiveness (R1), we use blue to indicate complete locking. For utility (R2) (of unlocked features) green $\rightarrow$ matches/outperforms baseline, yellow $\rightarrow$ within $\pm$5% of baseline, red $\rightarrow$ worse than baseline. Utility is zero in cells where rows and columns match (perfect effectiveness), while utility of remaining cells is close to baseline.
Locked Feature $\rightarrow$
Baseline
M + Q
M + S
M + U
Q + S
Q + U
S + U
M + Q + S
M + Q + U
M + S + U
Q + S + U
M + Q + S + U
DeepSeek-7B-Math locked via LOCKET
Math (M) 0.40 0.00 0.00 0.00 0.43 0.44 0.44 0.00 0.00 0.00 0.45 0.00
SQL (Q) 0.93 0.00 0.94 0.94 0.00 0.00 0.93 0.00 0.00 0.94 0.00 0.00
Summarize (S) 0.23 0.24 0.00 0.24 0.00 0.24 0.00 0.00 0.24 0.00 0.00 0.00
MMLU (U) 0.53 0.53 0.53 0.00 0.54 0.00 0.00 0.53 0.00 0.00 0.00 0.00
DeepSeek-7B-Coder locked via LOCKET
SQL (Q) 0.96 0.00 0.93 0.96 0.00 0.00 0.95 0.00 0.00 0.96 0.00 0.00
Llama-3-8B-Instruct locked via LOCKET
Math (M) 0.28 0.00 0.00 0.00 0.27 0.21 0.23 0.00 0.00 0.00 0.23 0.00
SQL (Q) 0.88 0.00 0.93 0.92 0.00 0.00 0.92 0.00 0.00 0.89 0.00 0.00
Summarize (S) 0.32 0.34 0.00 0.33 0.00 0.32 0.00 0.00 0.33 0.00 0.00 0.00
MMLU (U) 0.67 0.73 0.70 0.00 0.69 0.00 0.00 0.72 0.00 0.00 0.00 0.00
6.5 R4 (Scalable)
We evaluate scalability by measuring effectiveness (R1) and utility (R2) when locking multiple features. Table 4 shows results across all models and feature combinations, with color coding as in Table 2. LOCKET achieves perfect effectiveness (blue) with utility drops $\leq$7% from M-U interference.
Table 5 compares LOCKET and PWD with two or three features locked (M + Q, M + Q + S) for DeepSeek-7B-Math; other models show similar patterns (Appendix H). PWD’s utility and effectiveness are mostly worse (orange, yellow, or red), suggesting LOCKET scales better. PWD’s full fine-tuning likely causes “catastrophic forgetting” Kotha et al. (2024), where training refusal for one feature harms others.
Table 5: Comparison of LOCKET with prior work: Scalability w.r.t. R1 and R2 of LOCKET with prior work (“PWD”) Greenblatt et al. (2024a); Tang et al. (2024) locking DeepSeek-7B-Math (more in Appendix H). Color coding same as Table 4. For R1, we use blue to indicate complete locking and orange otherwise.
Locked Feature $\rightarrow$ M + Q M + Q + S
Eval. Feature $\downarrow$ PWD LOCKET PWD LOCKET
Math (M) 0.35 0.00 0.26 0.00
SQL (Q) 0.00 0.00 0.00 0.00
Summarize (S) 0.27 0.24 0.12 0.00
MMLU (U) 0.50 0.53 0.46 0.53
2
Takeaway: LOCKET outperforms PWD and meets all requirements (R1-R4).
7 Discussions and Summary
Overhead of LOCKET. We evaluate the runtime overhead of LOCKET on a single A100 40GB PCIe GPU, averaged over 5 runs. Attaching an adapter takes $1\pm 0.06$ second, while detaching takes $0.02\pm 0.00$ seconds. Attaching only needs to occur once per user session (i.e., at login), and inference latency and throughput remain unaffected: Time-to-First-Token (TTFT) is $3\pm 0.30$ms regardless of the number of adapters attached. These results show that serving costs for low-tier and paying users are comparable, making pay-to-unlock economically feasible. Storage size of individual LOCKET adapters is also a fraction of the base model size. A single locket adapter is only 1.6 - 1.7% of the base model’s total parameter count (120M for DeepSeek-7B-Math/DeepSeek-7B-Coder, 130M for Llama-3-8B-Instruct). Overall, LOCKET incurs minimal overhead in both storage and computation.
Model Types. Following prior work on password-locking Greenblatt et al. (2024a), we consider three model types (§6). To validate that our adapter-based feature-locking technique can scale effectively to larger models, we provide evaluation results on Llama-3-70B-Instruct in Appendix H. We found that LOCKET remains effective and utility-preserving (other than expected interference between M and U). We leave a more comprehensive evaluation across other models as future work.
Scalability Evaluation. We considered four features for the main evaluation, though our approach naturally extends to more. Prior work Lee et al. (2025) demonstrates that roughly 8 adapters can be merged with at most 15% utility drop. Our findings align with this trend: in Appendix I, we show for DeepSeek-7B-Math that LOCKET can scale to 8 features while keeping utility and effectiveness degradation within the 15% bound established in prior work. We leave further scalability improvements as future work.
Summary. We identify pay-to-unlock features as a new LLM application requiring FLoTEs that are effective, utility-preserving, robust, and scalable. No prior work meets all requirements. We propose LOCKET, the first robust and scalable FLoTE.
Limitations
  • Feature Interference. Ideally, features should be non-overlapping, but interference can occur in practice. We observed this in a few cases (Math and MMLU, which contains math-related questions); designing interference-resistant FLoTEs remains future work. Note that this interference is distinct from catastrophic forgetting, a failure mode of full fine-tuning methods like PWD where locking one feature overwrites knowledge of others. LOCKET does not suffer from forgetting because adapters attach dynamically to a frozen base model without permanently modifying its weights. We intend to explore potential mitigation strategies such as pre-processing feature datasets to remove semantically overlapping samples before adapter training in follow-up work.
  • Arms Race for Robustness. While we demonstrate robustness against state-of-the-art attacks, stronger future attacks may evade LOCKET. Since jailbreak attacks are relevant, defenses from that literature can be adopted, and LOCKET’s adapters can be fine-tuned accordingly.
  • Energy Consumption. While LOCKET complements existing tiered subscriptions, the energy consumption remains the same. Serving engines such as vLLM Kwon et al. (2023) could reduce costs further by efficiently loading and configuring adapters on a per-query basis.
Acknowledgments
This work is supported in part by Lambda (for cloud compute) and the Government of Ontario. Lipeng and Vasisht are supported by David R. Cheriton Graduate Scholarship. Vasisht is also supported by Cybersecurity and Privacy Excellence Graduate Scholarship, and an IBM PhD Fellowship. Views expressed in the paper are those of the authors and do not necessarily reflect the position of the funding agencies.
Appendix A Notations
Table 6: Frequently used notations and descriptions.
Notation Description
$\mathcal{F},f_{i},m$ Set of features, a feature, # of features.
$C,\mathcal{C}$ Client, Set of clients.
$a_{i}$ Adapter to lock (refuse) feature $f_{i}$.
$\pi_{\theta}$ Language model with parameters $\theta$.
$\mathcal{R}$ Responses generated by $\pi_{\theta}$
$l,L$ A layer index, and the set of target layers.
$D_{f_{i}}$ Dataset corresponding to feature $f_{i}$.
$(x_{i},y_{i})$ A prompt-response pair from $D_{f_{i}}$.
$D_{auth},D_{unauth}$ Datasets for utility and refusal training.
$c_{i},r_{i}$ Chosen and rejected responses.
$\mathcal{L}_{lock}$ Total loss for adapter fine-tuning.
$\mathcal{L}_{utility},\mathcal{L}_{robust}$ Losses for utility and robustness.
$\delta,\epsilon$ Perturbations and its L2-norm budget.
$\gamma(x,\delta)$ Function applying $\delta$ to activations.
$\Delta W^{i}$ Weight update matrix for adapter $a_{i}$.
$\sigma^{i},\sigma_{l}$ Adapter L2-norm; max norm over layer $l$.
$\tau$ Scaling hyperparameter for clipping.
$Clip_{l}=\tau\sigma_{l}$ Norm clipping threshold.
Appendix B Definition of a Feature
We define a feature as a specific model capability required to perform a task (e.g., “Math”, “Code”). The granularity of feature definitions is determined by the service provider based on their product offering; features can range from coarse (e.g., “Math”) to fine-grained (e.g., “Algebra”, “Geometry”).
Appendix C Inference-Time Control Flow
At inference time (Figure 1, Step ➌), the access control module identifies the set of features authorized for the requesting client based on their profile and subscription status. It then selects the adapters corresponding to all unauthorized features (Steps ➍–➎), merges them, and attaches the merged adapter to the frozen base model before inference (Step ➏). The adapters are trained on feature-specific datasets, causing the model to refuse queries that invoke unauthorized features while leaving authorized features unaffected.
Appendix D Other Applications of FLoTEs
Beyond pay-to-unlock schemes, FLoTEs have several promising potential applications. They can serve as an alignment alternative by locking dangerous features while preserving utility for legitimate tasks, providing a more granular and reversible approach than methods that modify base model weights. They can potentially enable more robust feature-level unlearning by locking access to specific features rather than attempting to erase them, making unauthorized knowledge more difficult to elicit even through adversarial prompting. They support staged feature releases, allowing providers to gradually roll out features to different user groups without maintaining multiple model versions. Finally, they enforce conditional compliance for sensitive features such as medical diagnoses or legal advice, keeping them locked by default and unlocking only for users or regions where appropriate regulatory approvals have been obtained.
Appendix E Implementation Details
Adapter Training. For LOCKET, we train LoRA adapters with a rank of $64$, alpha of $64$, and a dropout of $0.1$. We use RSLoRA Kalajdzievski (2023) for improved performance. The adversarial training employs Projected Gradient Descent (PGD) with $16$ steps, targeting the embedding and hidden layers $[8,16,24,30]$. We train for $100$ total steps with a batch size of $2$. For the baseline, we follow the SFT configurations of prior work Greenblatt et al. (2024b); Tang et al. (2024), using a validation set comprising $20\%$ no-password prompts and $80\%$ incorrect-password prompts to ensure robust refusal learning. For tuning the scaling threshold $\tau$ in our adapter merging strategy, we use a random sample of $100$ examples from each test set.
Dataset Composition. Train and test splits compositions of datasets can be found in Table 7. We use public open-sourced datasets and models.
Table 7: Datasets
Dataset Train Test
Training (utility) UltraChat 165,298 -
Feature (specific) SQL Create Context 62,861 15,716
MATH 7,500 5,000
Samsum 819 14,732
Feature (general) MMLU 99,842 14,042
Adapter Merging. We run hyperparameter tuning experiments to select optimal $\tau$ values for each feature combination. The following are the final $\tau$ values we use in merging the adapters for effectiveness and robustness evaluation, for DeepSeek-7B-Math, we have: M ($0.9$), Q ($0.7$), S ($0.5$), U ($0.7$), M + Q ($0.85$), M + S ($0.85$), M + U ($0.85$), Q + S ($0.6$), Q + U ($0.8$), M + Q + S ($0.75$), M + Q + U ($0.9$), M + S + U ($0.85$), Q + S + U ($0.75$), M + Q + S + U ($0.75$); for DeepSeek-7B-Coder, we have: Q ($0.45$); for Llama-3-8B-Instruct, we have M ($0.7$), Q ($0.6$), S ($0.9$), U ($0.8$), M + Q ($0.7$), M + S ($0.8$), M + U ($0.7$), Q + S ($0.8$), Q + U ($0.8$), M + Q + S ($0.8$), M + Q + U ($0.7$), M + S + U ($0.75$), Q + S + U ($0.8$), M + Q + S + U ($0.75$).
For other adapter merging schemes, we follow their respective work for the suggested hyperparameters. Specifically, for CAT Prabhakar et al. (2025) and Linear/Task Arithmetic Ilharco et al. (2023), we apply equal weights to each adapter during merging; for TIES Yadav et al. (2023), we use density = 0.5.
Jailbreak Attacks. We use a suite of state-of-the-art jailbreaking attacks to generate adversarial prompts, and evaluate robustness: (i) Many-shot JailbreakingAnil et al. (2024): For math, we sample $2$, 44, and $8$ in-context examples from problems below level-$2$ difficulty, and report the highest ASR achieved. For other tasks, we use randomly sampled demos. (ii) GCGZou et al. (2023): We run 125 iterations to optimize for the target string "Sure, here’s how to solve this problem," using a batch size of $64$ and replacing one character at a time. (iii) TAPMehrotra et al. (2024): We generate adversarial prompts using lmsys/vicuna-13b-v1.5-16k as the attack model, with a branching factor of $4$, width of $10$, depth of $5$, and the ground truth as the target. (iv) AutoDAN-TurboLiu et al. (2025): We run a single warm-up iteration with a size of $50$ for $150$ epochs, followed by one lifelong iteration. All attacks use $1,000$ random samples from each feature dataset, and generations are performed with a temperature of zero for deterministic outputs.
Computational Cost. Experiments are done using 8 * NVIDIA A100 40GB GPUs which consume in total round 6000 GPU hours.
Appendix F Variance Across Runs
Throughout our main experiments we set the inference temperature to zero (greedy decoding) to eliminate sampling variance, following prior work on password-locking Greenblatt et al. (2024b); Hofstätter et al. (2025) and jailbreak evaluation Mehrotra et al. (2024); Liu et al. (2025). To validate that our results hold under probabilistic sampling, we conducted 10 independent inference runs on DeepSeek-7B-Math with U locked, using temperature =1=1. Effectiveness remained near-perfect ($\leq\!1\%$ utility on unauthorized requests, $\pm 1\%$), and utility on the unlocked features was preserved ($\approx\!$ baseline $\pm 2\%$), confirming that the reported results are stable and not artefacts of greedy decoding.
Appendix G Adapter Merging
Table 8: Comparison of LOCKET merging with other commonly used merging techniques. LOCKET merging mitigates over-refusal while maintaining effective locking. ACC: Unlocked Feature Accuracy (higher is better), ALA: Averaged Locked Feature Accuracy (lower is better). All results are obtained on DeepSeek-7B-Math. LOCKET matches baseline accuracy on unlocked features while keeping locked features effectively refused (ALA==0). Other methods either over-refuse (ACC==0 on unlocked features) or fail to lock effectively (high ALA).
Unlocked Feature $\downarrow$ Baseline LOCKET Merging SVD TIES SVD DARE TIES DARE Linear DARE TIES SVD DARE Linear SVD Magnitude Prune Magnitude Prune SVD
Math ACC $\uparrow$ 0.40 0.45 0.00 0.00 0.40 0.00 0.00 0.00 0.00 0.00
ALA $\downarrow$ 0.00 0.00 0.00 0.00 0.37 0.02 0.00 0.02 0.01 0.01
SQL ACC $\uparrow$ 0.93 0.94 0.00 0.00 0.93 0.00 0.00 0.00 0.00 0.00
ALA $\downarrow$ 0.00 0.00 0.00 0.00 0.29 0.02 0.00 0.00 0.01 0.01
Summarize ACC $\uparrow$ 0.23 0.24 0.00 0.00 0.22 0.00 0.00 0.00 0.06 0.00
ALA $\downarrow$ 0.00 0.00 0.00 0.00 0.48 0.03 0.00 0.02 0.00 0.00
MMLU ACC $\uparrow$ 0.53 0.53 0.00 0.00 0.54 0.00 0.00 0.00 0.00 0.00
ALA $\downarrow$ 0.00 0.00 0.00 0.00 0.45 0.02 0.00 0.02 0.00 0.00
Matches baseline Over-refusing Over-refusing Ineffective locking Over-refusing Over-refusing Over-refusing Over-refusing Over-refusing
Table 8 compares LOCKET merging against standard adapter merging techniques avaialble in the PEFT library Mangrulkar et al. (2022). Most existing methods either over-refuse (refusing even unlocked features, ACC = 0) or fail to lock effectively (high ALA on locked features). LOCKET merging, however, can match baseline accuracy on unlocked features while maintaining perfect locking (ALA = 0).
Table 9: LOCKET remains reasonably scalable beyond four features: “Baseline” is the original model behaviour without FLoTE. For effectiveness (R1), we use blue to indicate complete locking, ineffective locking $\rightarrow$ orange. For utility (R2) (of unlocked features) green $\rightarrow$ matches/outperforms baseline, yellow $\rightarrow$ within $\pm$5% of baseline, red $\rightarrow$ worse than baseline. L, Y, P and O data are taken from the subsets of MMLU, which are largely within the same distribution, potentially causing larger feature interference and lead to less desirable results (as discussed in Limitations).
Locked Feature $\rightarrow$ Baseline M M + Q M + Q + S M + Q + S + L M + Q + S + L + H M + Q + S + L + H + Y M + Q + S + L + H + Y + P M + Q + S + L + H + Y + P + O
Math (M) 0.40 0.00 0.00 0.00 0.00 0.00 0.00 0.02 0.00
SQL (Q) 0.93 0.95 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Summarize (S) 0.23 0.23 0.24 0.00 0.00 0.00 0.00 0.00 0.00
Law (L) 0.35 0.35 0.35 0.35 0.00 0.00 0.09 0.10 0.00
History (H) 0.54 0.56 0.52 0.52 0.46 0.02 0.12 0.09 0.00
Psychology (Y) 0.64 0.63 0.63 0.62 0.56 0.54 0.08 0.11 0.00
Politics (P) 0.44 0.44 0.43 0.43 0.39 0.45 0.38 0.06 0.00
Philosophy (O) 0.49 0.45 0.45 0.44 0.43 0.40 0.44 0.34 0.00
Appendix H Additional Scalability Results
We provide additional comparisons between LOCKET and prior work across different models and feature combinations. The first two tables (Table 10, 11) compare LOCKET with password-locking (PWD) on Llama-3-8B-Instruct and DeepSeek-7B-Coder, showing that LOCKET consistently outperforms PWD in preserving utility while maintaining effective locking. The third table (Table 12) compares LOCKET against PWD with Circuit Breaking (PWD + CB) Hofstätter et al. (2025) on DeepSeek-7B-Math; while PWD+CB improves robustness, it fails to maintain effectiveness when locking multiple features and still degrades utility on U, whereas LOCKET achieves perfect effectiveness across all configurations. The fourth table (Table 13) demonstrates that LOCKET scales to larger models (Llama-3-70B-Instruct, 4-bit quantized), remaining effective and utility-preserving with only minor U degradation due to interference with M.
Table 10: Comparison of LOCKET with prior work: Scalability w.r.t. R1 and R2 of LOCKET with prior work (“PWD”) Greenblatt et al. (2024a); Tang et al. (2024) locking Llama-3-8B-Instruct. Color coding for scalability, are same as Table 5.
Locked Feat. $\rightarrow$ M M + Q M + Q + S
Eval. Feat. $\downarrow$ PWD LOCKET PWD LOCKET PWD LOCKET
Math (M) 0.00 0.00 0.00 0.00 0.00 0.00
SQL (Q) 0.01 0.92 0.00 0.00 0.00 0.00
Summarize (S) 0.26 0.34 0.41 0.34 0.51 0.00
MMLU (U) 0.06 0.64 0.05 0.73 0.05 0.72
Table 11: Comparison of LOCKET with prior work: Scalability w.r.t. R1 and R2 of LOCKET with prior work (“PWD”) Greenblatt et al. (2024a); Tang et al. (2024) locking DeepSeek-7B-Coder. Color coding for scalability, are same as Table 5.
Locked Feat. $\rightarrow$ M M + Q M + Q + S
Eval. Feat. $\downarrow$ PWD LOCKET PWD LOCKET PWD LOCKET
SQL (Q) 0.01 0.96 0.00 0.00 0.00 0.00
Table 12: Comparison of LOCKET with prior work: Scalability w.r.t. R1 and R2 of LOCKET with prior work (“PWD + CB”) Hofstätter et al. (2025) locking DeepSeek-7B-Math via PWD and Circuit Breaking. Color coding for scalability, are same as Table 5.
Locked Feat. $\rightarrow$ M M + Q M + Q + S
Eval. Feat. $\downarrow$ P.+CB LOCKET P.+CB LOCKET P.+CB LOCKET
Math (M) 0.00 0.00 0.38 0.00 0.20 0.00
SQL (Q) 0.93 0.95 0.47 0.00 0.05 0.00
Summarize (S) 0.23 0.23 0.28 0.24 0.50 0.00
MMLU (U) 0.38 0.51 0.43 0.53 0.39 0.53
Table 13: Applying LOCKET to a larger-scale model: Scalability w.r.t. R1 and R2 of LOCKET locking Llama-3-70B-Instruct (4-bit quantized). Color coding for scalability, are same as Table 5.
Locked Feat. $\rightarrow$ None M M + Q M + Q + S
Math (M) 0.49 0.00 0.00 0.00
SQL (Q) 0.98 0.98 0.00 0.00
Summarize (S) 0.37 0.38 0.37 0.00
MMLU (U) 0.81 0.79 0.78 0.75
Appendix I Scalability beyond Four Features
We evaluate LOCKET’s scalability beyond four features by progressively locking up to eight features on DeepSeek-7B-Math (Table 9). The additional features (Law, History, Psychology, Politics, Philosophy) are drawn from MMLU subcategories, which share similar distributions and thus exhibit greater feature interference. Despite this challenging setting, LOCKET maintains perfect effectiveness in most configurations, with only minor ineffective locking (orange cells) appearing when six or more features are locked simultaneously. Utility degradation increases with the number of locked features, particularly among MMLU subcategories due to their distributional overlap. These results align with prior work Lee et al. (2025) showing that adapter merging can scale to 8 adapters with roughly 15% degradation, confirming that LOCKET remains reasonably scalable beyond the four features evaluated in the main text.